# Multi-Exit Kolmogorov–Arnold Networks: enhancing accuracy and parsimony

### v1

@@ 4.3 Continual learning
[…]
@@ 4.4 Real-world data
  • Power Plant Energy. Predict the electrical power output (in MW) for different ambient atmospheric conditions, temperature, pressure, relative humidity, and the steam turbine pressure (or vacuum). Data originated from a 480 MW combined cycle power plant with two gas turbines, one steam turbine, and two heat recovery steam generators, collected over a six-year period (2006-2011) [49, 50].
@@
[…]
@@ 5 Learning to exit
The exit weight hyperparameter complicates the architecture search problem, already a potential challenge when estimating KAN models. Indeed, from our experiments, it is not always obvious a priori what exit weights will best optimize the KAN. Sometimes uniform weights work well, or an increasing or decreasing ramp, or sometimes even a heavy weight on the last exit(s) can be beneficial. While KANs train quite quickly on these smaller scientific problems, allowing rapid iteration to explore the weight simplex, nevertheless, it would be useful, and data-efficient, to avoid this trial-and-error process. To this end, here we introduce and apply a basic “learning to exit” algorithm which treats the exit weights as learnable parameters, eliminating them entirely from the architecture search.
@@ 5.1 Learnable exit weights
where $w_{i}$ ($i=0,\ldots,L-1$) are predetermined constants and $\mathcal{L}_{i}$ is the loss function for exit $i$. In our learning to exit framework, we introduce exit logits $\bm{\theta}_{w}=\{\theta_{0},\theta_{1},\ldots,\theta_{L-1}\}$ to be optimized. A softmax transformation connects these to Eq. 17,
@@ 5.3 Results
[…]
@@
The dynamics of exit selection are worth further study—for one, we expected a regularization term on the exit logits would be necessary, but these results imply otherwise—yet our findings already show that the learning-to-exit model has promise.
@@ 6 Discussion
Despite the promising results presented in this work, some limitations should be acknowledged. Perhaps the most serious concern is the extra need to set the exit weights when fitting a multi-exit KAN. In all our experiments, setting $w$ required only brief coarse-grained tuning (and it is addressed by our learning-to-exit algorithm) but in settings where data are scarce, it may be difficult to optimize the exit weights without overfitting. Second, we focused our comparisons on single-exit versus multi-exit KANs to isolate the effect of the architectural change, leaving broader comparisons for future work. Likewise, future work should consider the effects of different hyperparameter values and training settings, including other sources of regularization such as pruning [15], since it is important in practice to ascertain the optimal settings for different applications and whether multi-exits benefit from or remain useful with such fine-tuning. Third, while multi-exit architectures often identify more parsimonious models, the interpretability gains are indirect—the exits themselves do not enhance the interpretability of individual activation functions, but rather help identify simpler network configurations. Fourth, our learning-to-exit algorithm should be studied further and, while effective, can surely be improved. Finally, the additional computational overhead during training, though modest, may become more significant for very deep networks with many exits.
@@
Several promising directions are worth pursuing. First, receiving multiple predictions from a KAN immediately brings to mind the idea of ensemble learning [54]. However, ensembles benefit from uncorrelated or de-correlated models, but the different exits in a multi-exit KANs are not independent. Investigating methods to encourage heterogeneity among exits while maintaining their collaborative training could enable ensemble learning. This may also lead to uncertainty quantification capabilities [55, 56]—a key goal for KANs [57, 58]——through the natural variation of predictions across exits. Second, our learning-to-exit framework leaves room for improvement, and incorporating ideas from differentiable architecture search [59] more generally could benefit KANs. Third, exploring exit architectures beyond simple single-layer KAN exits may yield better performance, although doing so without harming parsimony may be difficult. Fourth, extending multi-exit architectures from B-splines to other KAN variants (Fourier-KANs, Wavelet-KANs, Physics-Informed KANs, etc.) could reveal whether the benefits generalize across different basis functions. Finally, applying multi-exit KANs to larger-scale scientific problems, particularly in domains like climate modeling or molecular dynamics where both accuracy and interpretability are crucial, would provide valuable insights into their practical utility.
@@ Conclusion
We have introduced multi-exit architectures for Kolmogorov–Arnold Networks, demonstrating that augmenting KANs with additional outputs consistently improves their performance across diverse scientific modeling tasks. Our experiments revealed that multi-exit KANs often achieve their best performance at earlier exits, indicating that they successfully identify more parsimonious models without sacrificing—and often improving—accuracy. While multi-exit architectures introduce additional hyperparameters, our results show the benefits substantially outweigh this added complexity This finding is particularly valuable for scientific applications where interpretability is paramount, as simpler models with fewer compositional layers are inherently more interpretable. These results suggest that multi-exit architectures represent a natural and effective enhancement to KANs, offering researchers a principled approach to finding the right balance between model complexity and performance.
@@ Appendix A Materials and Methods
KANs were implemented with the PyKAN library v0.2.8 (https://github.com/KindXiaoming/pykan), based on PyTorch v2.6.0 [60]. To fit KAN models using training data, the Limited-memory Broyden-Fletcher-Goldfarb-Shanno (L-BFGS) algorithm [35] was used with history size 10, strong Wolfe conditions for line search, and convergence tolerances of $10^{-32}$ for gradient norm, parameter changes, and curvature conditions. Unless otherwise noted, fitting employed a progressive grid refinement strategy, iteratively refining the spline basis functions through a sequence of increasing grid sizes $G=3,5,10,20$, with 30 optimization steps performed at each grid resolution. This approach allows the model to first capture coarse-grained patterns before learning finer details, leading to more stable convergence and better generalization performance [15]. All other training settings and hyperparameters were kept at the default values of PyKAN, including the learning rate of 1.0 and the default spline order $K=3$ and grid update schedule. Multiplication units were not used. The same training procedure and settings was used throughout, for both single-exit and multi-exit KANs. Fine-tuning these settings would likely further improve performance, but both types of KANs would be expected to benefit. Our source code will be available at https://github.com/bagrow/multi-exit-KAN upon acceptance of this manuscript.
@@
For the 1D and 2D regression tasks, $n=1000$ training and $n=1000$ testing points were generated, with $\mathbf{x}\sim U\left([x_{\min},x_{\max}]^{d}\right)$ and $y$ generated from $\mathbf{x}$, without additional noise, according to the given equation. For the fits illustrated in Fig. 1, $n=1000$ training points and $n=200$ testing points were used, as well as a single $G=5$ grid, 30 optimization steps, KAN shape $[2,3,2,1]$ and $w=[1,1,1]$. The Feynman Equations dataset used the same data generating process as the 1D and 2D regression tasks, but each exogenous variable’s range was given by the range of values in the released AI Feynman dataset [42]. For the dynamical systems experiments, data for each system was generated and split into training and testing folds following the procedure and parameters of Panahi et al. [19]. The continual learning experiment used 100 samples per peak, a grid size of 20 without refinement or updates, 10 L-BFGS steps per phase, and all other settings at PyKAN defaults. Details for the three real-world datasets were covered in Sec. 4.4.

### v2

@@ 4.3 Continual learning
As seen in Fig. 6, we can answer in the affirmative. While both architectures display some forgetting, with previously learned peaks changing with subsequent data, the effect is worse for the single-exit KAN (shape $[1,5,5,1]$), especially when learning the second peak. Compared to the single-exit KAN of the same architecture (Fig 6, middle row), the multi-exit KAN (bottom row) with exit weights $w=[1,1,2\text{e}3]$ better tracks the underlying function across training phases, and when finished displays less than half the error of the single-exit KAN ($\mathit{RMSE}=0.086$ vs. $0.19$).
@@ 4.4 Real-world data
  • Power Plant Energy. Predict the electrical power output (in MW) for different ambient atmospheric conditions, temperature, pressure, relative humidity, and the steam turbine pressure (or vacuum). Data originated from a 480 MW combined cycle power plant with two gas turbines, one steam turbine, and two heat recovery steam generators, and were collected over a six-year period (2006-2011) [49, 50].
@@
As seen in Table 2, multi-exits improved predictive performance on all three datasets. Interestingly, for all datasets, multi-exit KANs outperformed at early exits the single-exit KANs of the same size and smaller. For instance, on Airfoil, the three-exit KAN achieved $\mathit{RMSE}=5.129$ at Exit 0, improving on the single-exit KAN of the same size as that exit ($\mathit{RMSE}=5.338$) by 3.9%. To assess generalizability of these improvements, we evaluated performance on additional test data for two datasets (omitting Airfoil due to data availability). A one-sided Wilcoxon signed-rank test compared squared residuals between the best single-exit network and the best multi-exit KAN at its optimal exit (as identified in Table 2). In both datasets, the residuals of the single-exit network were significantly larger than those of the multi-exit KAN (Power plant: W=15 968 151W=15\,968\,151, $p<0.001$, n=7 568n=7\,568; Superconductivity: W=25 745 299W=25\,745\,299, $p<0.01$, n=10 000n=10\,000). These results confirm that the multi-exit architecture provides statistically significant improvements in predictive accuracy that generalize to new unseen data.
Multi-exit KANs achieved accurate, generalizable fits while simultaneously being more parsimonious on all three datasets. While additional training of the single-exit KANs could potentially close these gaps, these results nevertheless suggest that multi-exit KANs improve performance simultaneously across shallower and deeper architectures.
@@ 5 Learning to exit
The exit weight hyperparameter complicates the architecture search problem, already a potential challenge when estimating KAN models. Indeed, from our experiments, it is not always obvious a priori what exit weights will best optimize the KAN. Sometimes uniform weights work well, or an increasing or decreasing ramp, or sometimes even a heavy weight on the last exit(s) can be beneficial. While KANs train quite quickly on these smaller scientific problems, allowing rapid iteration to explore the weight simplex, nevertheless, it would be useful, and data-efficient, to avoid this trial-and-error process. To this end, here we introduce and apply a basic “learning-to-exit” algorithm which treats the exit weights as learnable parameters, eliminating them entirely from the architecture search.
@@ 5.1 Learnable exit weights
where $w_{i}$ ($i=0,\ldots,L-1$) are predetermined constants and $\mathcal{L}_{i}$ is the loss function for exit $i$. In our learning-to-exit framework, we introduce exit logits $\bm{\theta}_{w}=\{\theta_{0},\theta_{1},\ldots,\theta_{L-1}\}$ to be optimized. A softmax transformation connects these to Eq. 17,
@@ 5.3 Results
To evaluate the learning-to-exit model, we first apply it to the three datasets studied in Sec. 4.4, focusing on the three-exit architectures. Comparing to Table 2, the results were promising. On every dataset, the learned model outperformed the single-exit KAN of the same shape. Further, in two of the three cases, the new model outperformed every model in Table 2. Specifically, for Airfoil the new model achieved $\mathit{RMSE}=4.947$ and for Superconductivity $\mathit{RMSE}=18.126$, beating the previous best $\mathit{RMSE}=5.129$ and $18.353$, respectively. On the other hand, for Power plant, it achieved $\mathit{RMSE}=4.558$, better than the single-exit result of $\mathit{RMSE}=4.601$ but worse than the multi-exit’s $\mathit{RMSE}=4.406$.
@@
Next, we consider the method on some of the earlier synthetic nonlinear regression tasks. For the sinc function (Fig. 2), the learning-to-exit algorithm with uniform initial weight converged to weights $w=[0.001,0.003,0.761,0.236]$, favoring Exit 2. This configuration achieved $\mathit{RMSE}=0.00170$, an 89% improvement over the single-exit baseline ($\mathit{RMSE}=0.0154$) although not outperforming the fixed weight result of $\mathit{RMSE}=0.00145$ (which used the deeper Exit 3). For the 2D regression (Fig. 3), using a decreasing weight initialization (logits $\bm{\theta}_{w}=[1,0,-1]$), the algorithm converged to weights $w\approx[0,1,0]$, focusing entirely on Exit 1 with $\mathit{RMSE}=0.00756$, a 67% improvement over the single-exit baseline ($\mathit{RMSE}=0.0232$) though slightly worse than our previous best results of $\mathit{RMSE}=0.0045$ and $0.0071$. A decreasing weight initialization outperformed a uniform one for this task, biasing toward earlier exits and promoting parsimony, suggesting that initialization strategies merit further investigation.
Finally, for the dynamical systems, results were mixed. On the food chain ecosystem, the learning-to-exit algorithm with increasing weight initialization converged to $w=[0.108,0.113,0.779]$ but achieved $\mathit{RMSE}=5.18\times 10^{-4}$, underperforming both the single-exit baseline ($\mathit{RMSE}=3.28\times 10^{-4}$) and the previous fixed-weight result ($\mathit{RMSE}=3.77\times 10^{-4}$). For the Ikeda map, the algorithm with uniform initialization converged to $w\approx[0,1,0,0]$, focusing on Exit 1 with $\mathit{RMSE}=4.66\times 10^{-3}$, not matching the previous fixed-weight performance ($\mathit{RMSE}=4.56\times 10^{-3}$) but showing a modest improvement over the single-exit baseline ($\mathit{RMSE}=4.82\times 10^{-3}$).
Overall, our results highlight the learning-to-exit approach while revealing that performance depends critically on initialization strategies and problem characteristics. The dynamics of exit selection warrant further study—for one, we expected a regularization term on the exit logits would be necessary, but these results suggest otherwise—yet our findings already show that the learning-to-exit model has promise.
@@ 6 Discussion
Multi-exit models outperformed all same size and smaller KANs on the real world data (Table 2), yet we may encounter cases where the best model comes from directly training the appropriate size KAN without using multi-exits. Indeed, we know this happens, for instance in the sinc function results (Fig. 2). A single-exit KAN with one fewer layer than we presented achieved excellent performance ($\mathit{RMSE}=1.91\times$10-4), suggesting that the exit weights shown in the figure were suboptimal. This observation reveals a key insight: every shallower single-exit network (with otherwise the same layer widths) represents a special case of the multi-exit architecture, corresponding to one-hot exit weights $w=[1,0,\ldots],[0,1,0,\ldots],\ldots,[0,\ldots,0,1]$—the vertices of the exit weight simplex. While these vertices may represent good solutions, multi-exit architectures can relax into the interior of the simplex to find even better performance. Indeed, optimizing the sinc function’s exit weights to $w=[2.5\times 10^{-4},2.2\times 10^{-2},1\times 10^{4},0]$ (unnormalized), an interior solution near the vertex corresponding to the better-performing shallower network, improved performance to $\mathit{RMSE}=1.80\times 10^{-4}$ at the third exit—a relative improvement of over 5%. This example demonstrates that multi-exit architectures provide both architecture search capabilities and the potential to discover superior solutions through continuous weight optimization.
While multi-exit KANs introduce additional complexity through exit weights, this does not compromise KAN’s core interpretability advantages. The activation functions remain visualizable and amenable to symbolic regression [16], while learned exit weights can themselves provide interpretable insights into layer contributions during training. Most importantly, multi-exit KANs improve interpretability by identifying more parsimonious configurations—when performance is achieved at early exits, fewer layers of functional composition are needed, yielding inherently more interpretable models.
@@
Despite the promising results presented in this work, some limitations should be acknowledged. Perhaps the most serious concern is the extra need to set the exit weights when fitting a multi-exit KAN. In all our experiments, setting $w$ required only brief coarse-grained tuning (and it is addressed by our learning-to-exit algorithm) but in settings where data are scarce, it may be difficult to optimize the exit weights without overfitting. Second, we focused our comparisons on single-exit versus multi-exit KANs to isolate the effect of the architectural change, leaving broader comparisons for future work. While comparisons of single- and multi-exit KANs with other interpretable architectures such as decision trees, linear models, or attention-based interpretable networks would provide valuable context, comparing methods with different notions of interpretability calls for different evaluation protocols. Such systematic comparisons across different interpretable architectures are beyond the scope of this work but represent an important direction for future research. Likewise, future work should consider the effects of different hyperparameter values and training settings, including other sources of regularization such as pruning [15], to ascertain the optimal settings for different applications and whether multi-exits benefit from or remain robust to such techniques. Pruning, in particular, while non-differentiable, may interplay with the different exit layers in interesting and useful ways. Third, while multi-exit architectures often identify more parsimonious models, the interpretability gains are indirect—the exits themselves do not enhance the interpretability of individual activation functions, but rather help identify simpler network configurations. Fourth, our learning-to-exit algorithm should be studied further and, while effective, can surely be improved. Finally, the additional computational overhead during training, though modest, may become more significant for very deep networks with many exits.
@@
Several promising directions are worth pursuing. First, receiving multiple predictions from a KAN immediately brings to mind the idea of ensemble learning [54]. However, ensembles benefit from uncorrelated or de-correlated models, but the different exits in a multi-exit KANs are not independent. Investigating methods to encourage heterogeneity among exits while maintaining their collaborative training could enable ensemble learning. This may also lead to uncertainty quantification capabilities [55, 56]—a key goal for KANs [57, 58]——through the natural variation of predictions across exits. Second, our learning-to-exit framework leaves room for improvement, and incorporating ideas from differentiable architecture search [59] more generally could benefit KANs. Third, exploring exit architectures beyond simple single-layer KAN exits may yield better performance, although doing so without harming parsimony may be difficult. Fourth, extending multi-exit architectures from B-splines to other KAN variants (Fourier-KANs, Wavelet-KANs, Physics-Informed KANs, etc.) could reveal whether the benefits generalize across different basis functions. (We address this in part in App. B, but more should be done.) Finally, applying multi-exit KANs to larger-scale scientific problems, particularly in domains like climate modeling or molecular dynamics where both accuracy and interpretability are crucial, would provide valuable insights into their practical utility.
@@ Conclusion
We have introduced multi-exit architectures for Kolmogorov–Arnold Networks, demonstrating that augmenting KANs with additional outputs consistently improves their performance across diverse scientific modeling tasks. Our experiments revealed that multi-exit KANs often achieve their best performance at earlier exits, indicating that they successfully identify more parsimonious models without sacrificing—and often improving—accuracy. While multi-exit architectures introduce additional hyperparameters, our results show the benefits substantially outweigh this added complexity. This finding is particularly valuable for scientific applications where interpretability is paramount, as simpler models with fewer compositional layers are inherently more interpretable. These results suggest that multi-exit architectures represent a natural and effective enhancement to KANs, offering researchers a principled approach to finding the right balance between model complexity and performance.
@@ Appendix A Materials and Methods
KANs were implemented with the PyKAN library v0.2.8 (https://github.com/KindXiaoming/pykan), based on PyTorch v2.6.0 [60]. To fit KAN models using training data, the Limited-memory Broyden-Fletcher-Goldfarb-Shanno (L-BFGS) algorithm [35] was used with history size 10, strong Wolfe conditions for line search, and convergence tolerances of $10^{-32}$ for gradient norm, parameter changes, and curvature conditions. Unless otherwise noted, fitting employed a progressive grid refinement strategy, iteratively refining the spline basis functions through a sequence of increasing grid sizes $G=3,5,10,20$, with 30 optimization steps performed at each grid resolution. Spline regularization was not used ($\lambda=0$) as we found for our data and training settings that it always lowered performance. Instead, grid refinement acts as implicit regularization and allows the model to first capture coarse-grained patterns before learning finer details, leading to more stable convergence and better generalization performance [15]. Unless otherwise noted, all other training settings and hyperparameters were kept at the default values of PyKAN, including the learning rate of 1.0 and the default spline order $K=3$ and grid update schedule (grid points were equally spaced and updated every 10 optimization steps). Multiplication units were not used. The same training procedure and settings were always used for corresponding single-exit and multi-exit KANs. Fine-tuning these settings would likely further improve performance, but both types of KANs would be expected to benefit. Training times were modest with these hyperparameters and data, generally taking 30–90 s for single-exit KANs on a MacBook Pro (M1 Max CPU); the one exception being training on the Ikeda map that took approximately 7 min due in part to the dataset’s high sampling rate. Multi-exit KANs generally require 50–80% more time to train than the corresponding single-exit KAN, in line with our parameter count estimates (Sec. 3). Our source code is available at https://github.com/bagrow/multi-exit-KAN.
@@
For the 1D and 2D regression tasks, $n=1000$ training and $n=1000$ testing points were generated, with $\mathbf{x}\sim U\left([x_{\min},x_{\max}]^{d}\right)$ and $y$ generated from $\mathbf{x}$, without additional noise, according to the given equation. For the fits illustrated in Fig. 1, $n=1000$ training points and $n=200$ testing points were used, as well as a single $G=5$ grid, 30 optimization steps, KAN shape $[2,3,2,1]$ and $w=[1,1,1]$. The Feynman Equations dataset used the same data generating process as the 1D and 2D regression tasks, but each exogenous variable’s range was given by the range of values in the released AI Feynman dataset [42]. For the dynamical systems experiments, data for each system was generated and split into training and testing folds following the procedure and parameters of Panahi et al. [19]. For training, a learning rate of 0.1 was used for both systems, and for the Ikeda map specifically, grid updates were not used and 50 optimization steps per grid resolution instead of 30 were used. Other training settings were unchanged from other experiments. The continual learning experiment used 100 samples per peak, a grid size of 20 without refinement or updates, 10 L-BFGS steps per phase, and all other settings at PyKAN defaults. Details for the three real-world datasets were covered in Sec. 4.4.
Appendix B Multi-Exit Fourier KANs
To demonstrate that the benefits of multi-exits are not specific to the B-spline formulation, we implemented a Fourier-based variant as preliminary validation of multi-exits across different activation function representations.
Fourier KAN formulation
Following [61] (see also [32]), each activation function $\phi_{\ell,j,i}(x)$ connecting unit $i$ in layer $\ell$ to unit $j$ in layer $\ell+1$ is parameterized using Fourier series:
$\phi_{\ell,j,i}(x)=\sum_{k=1}^{G}\left[a_{\ell,j,i,k}\cos(\omega_{k}x)+b_{\ell,j,i,k}\sin(\omega_{k}x)\right]$ (23)
where $a_{\ell,j,i,k}$ and $b_{\ell,j,i,k}$ are learnable coefficients, $G$ is the number of frequency components (analogous to grid size in B-spline KANs), and $\omega_{k}=k\omega_{0}$ are the fundamental frequencies scaled by a factor $\omega_{0}$ to prevent high-frequency oscillations over the input domain. Coefficients are initialized from a normal distribution $\mathcal{N}(0,\sigma^{2})$ with $\sigma=1/\left(\sqrt{d}\left(k+1\right)^{2}\right)$ to bias learning toward lower frequencies.
The layer output aggregates contributions from all input connections:
$x_{\ell+1,j}=\sum_{i=1}^{N_{\ell}}\sum_{k=1}^{G}\left[a_{\ell,j,i,k}\cos(\omega_{k}x_{\ell,i})+b_{\ell,j,i,k}\sin(\omega_{k}x_{\ell,i})\right]+b_{j}$ (24)
where $b_{j}$ is a learnable bias term.
Multi-exits are added and trained as before (Sec. 3) with the exception that analogues of grid refinement and grid updates are not used.
Experimental validation
We tested Fourier KANs on the polynomial $f(x)=x^{3}-2x^{2}+x+1$ over $x\in[0,3]$, deliberately choosing a non-trigonometric target to avoid favoring the Fourier parameterization. This parallels our use of the sinc function (Eq. (12)) for B-spline KANs, ensuring neither method receives an unfair advantage.
We employed a shape $[1,2,2,1]$ architecture with $G=5$ frequency components and $\omega_{0}=\pi/10$, trained using L-BFGS optimization for 30 steps.
The single-exit Fourier KAN achieved an $\mathit{RMSE}=0.00230$. Multi-exit variants with two exit weight configurations demonstrated the effectiveness of the approach:
  • $w=[0,1,100]$: The final exit (Exit 2) achieved $\mathit{RMSE}=0.00145$, a 37% improvement.
  • $w=[0,100,1]$: The intermediate exit (Exit 1) achieved $\mathit{RMSE}=0.000761$, a 67% improvement. This configuration also outperformed a shape [1,2,1] (same as Exit 1) single-exit KAN that achieved $\mathit{RMSE}=0.00113$.
Notably, the second configuration shows that multi-exit training can identify more parsimonious solutions: the intermediate exit achieved superior performance compared to both the single-exit baseline and the deeper final exit, indicating that a shallower network architecture was sufficient for this task.
These results confirm that multi-exit architectures generalize beyond B-spline parameterizations, improving performance and identifying optimal depths across multiple KAN variants.